# Large Language Models, scientific knowledge and factuality: A systematic analysis in antibiotic discovery

### v2

@@ 3.1 Dataset
The investigation of the biological knowledge stored in LLMs requires a collection of biological relations that are elicited and referred in the scientific literature. Scientific publications are part of the training corpus for most LLMs, thus higher prevalence in literature should lead to better performance in factual knowledge extraction. In our study we created a dataset of biological relations accounting for the amount of supporting evidence (number of associated papers).
Chemical compound-fungus pairs were extracted from the Wikidata Knowledge Graph using the predicate found in taxon (p:P703), along with the number of bibliographic references supporting each relationship (via API). Classification and nomenclature information on fungi were extracted from the Mycobank database [^2] https://www.mycobank.org. Information about antibiotic activity of chemicals were extracted from the ChEBI ontology [^3] https://www.ebi.ac.uk/chebi/downloadsForward.do.
A subset containing only fungal species producing chemicals with antibiotic activity was first created. The pairs with the most bibliographic references were progressively integrated until obtaining 100 distinct subject fungi, which corresponds to 123 pairs. Then, a complementary subset (with an equal number of pairs) involving produced chemicals but without antibiotic activity was added from the same list of fungi. Again, the pairs with the most bibliographic references were selected. Only molecules with a 3-STAR review in ChEBI were considered and only those that have an annotated role in the antimicrobial agent class (CHEBI:33281) were selected for positive examples of antibiotics.
As a result, a dataset with 246 chemical compound-fungus pairs (see Figure 2, N) equally divided between chemicals with and without antibiotic activity was obtained. For the experiments, we randomly selected ten chemicals distributed by antibiotic activity and number of references (five chemicals with and five without antibiotic activity) (see Table S.1). These ten chemicals paired with one or more fungi. In total, there were 23 chemical compound-fungus pairs (see Figure 2, $n_{relations}$) in the analyzed subset. The selected ten chemical compounds were defined as entity 1 for two tasks. The fungus was defined as entity 2 in the second task. Of note, the selected 23 chemical-fungus relations are described by multiple scientific publications in PubMed, all published before 2019 (for exact numbers see Table S.1). This ensures these relations are contained in the LLM training corpus, provided that the corpus contains PubMed articles or abstracts.
@@ 3.2 Large Language Models
Table 1: Summary of the evaluated Large Language Models ordered by the release date.
Model Lab Training Corpus Vocabulary size Model size (parameters) used/available Date of the model release* Reference
GPT-2 OpenAI WebText: 40 GB of text, 8M documents, from 45M webpages upvoted on Reddit (all Wikipedia pages removed) 50,257 124M/1.5B Feb 2019 [6]
GPT-3 OpenAI 570 GB plaintext, 0.4 trillion tokens. (410B CommonCrawl, 19B WebText2, 3B English Wikipedia, and two books corpora (12B Books1 and 55B Books2)) 175B June 2020 [7]
GPT-neo EleutherAI The Pile - a large scale curated dataset created by EleutherAI 825GB, 380 billion tokens 50,257 1.3B/2.7B March 2021 [24]
Bloom BigScience 45 natural languages 12 programming languages In 1.5TB of pre-processed text, converted into 350B unique tokens 250,680 560M-3B/176B July 2022 [25]
BioGPT Microsoft 15M PubMed abstracts (trained from scratch) 42,384 347M/347M Oct 2022 [26]
BioGPT-large Microsoft 15M PubMed abstracts (trained from scratch) 42,384 1.5B/1.5B Oct 2022 [26]
Galactica Meta AI 106B tokens; Documents: 48M papers, 2M code, 8M reference material, 2M knowledge bases, 0.9M filtered commoncrawl, 1.3M prompts, 0.02M other 50,000 1.3B/120B Nov 2022 [27]
ChatGPT OpenAI unavailable 20B (175B) Nov 2022
GPT-4 OpenAI unavailable Feb 2023
7B
Llama 2 Meta AI unavailable 13B July 2023 [19]
70B
@@
In this study, we evaluated ten LLMs (summarized in Table 1), where six of them were released since July 2022. The models vary in terms of training corpus, vocabulary size, number of parameters, layers and maximum input sequence length. Below, we articulate the essential properties of each target model.
There are two main high-level architectures associated with language models: BERT - Bidirectional Encoder Representations from Transformers (and its derivatives) and GPT - ‘Generative Pre-trained Transformer’ (and its derivatives). Extensive research has been conducted on the former in the biomedical domain, demonstrating considerable success in numerous discriminative downstream tasks [6, 7]. In this paper, we evaluate generative transformer language models pre-trained on large scale general literature (GPT-2, GPT-3, GPT-neo, Galacitca, Bloom, ChatGPT, GPT-4, Llama 2) and specific for the biomedical domain (BioGPT and BioGPT-Large). Due to evaluation in the zero-shot setting, the study focuses on decoder-only architectures due to their superior performance in comparison to encoder-decoder frameworks [28].
GPT is a language generation model that utilizes the transformer decoder architecture for pre-training on large-scale text corpora. GPT is grounded in the next-token prediction task, in which the model learns to predict the next word token based on the preceding word tokens.
GPT-2 adopts GPT-1’s architecture with some differences, including the modified initialization, pre-normalization, and reversible tokenization. A different initialization, accounting for residual path accumulation, scales weights by 1/N, where N is the number of residual layers. Vocabulary expands to 50,257, context size increases to 1024 tokens, and batch size grows to 512. GPT-2 used a larger architecture and a more diverse corpus than GPT-1, including web pages, books, and scientific articles. GPT-3 uses the same model and architecture as GPT-2, with the exception of the use of alternating dense and sparse attention patterns in the layers of the transformer, similar to the Sparse Transformer. Compared to previous generative models, GPT-3 has an order of magnitude larger parameter space (175 billion parameters).
GPT-Neo is an autoregressive transformer language model, which is based on the replication of the GPT-3 architecture. With 2.7 billion parameters, the model was trained on the Pile [24, 29], a large-scale curated corpus containing Pubmed Central and Wikipedia data. The architecture of Bloom (BigScience Large Open-science Open-access Multilingual Language Model [25]) is similar to GPT-3, being trained on 46 different languages and 13 programming languages, including the Pile. The largest Bloom contains 176B parameters, however due to hardware access limitations, we tested Bloom under the 560M and 3B parameter setting.
BioGPT and BioGPT-large are domain-specific generative transformer language models pre-trained on large-scale biomedical literature [30], and are pre-trained on 15M PubMed items (each with both title and abstract) from scratch. We tested both versions: with 347M and 1.5B parameters (BioGPT-large). BioGPT uses the GPT-2 architecture as its backbone. However, instead of relying on the GPT-2 vocabulary, it has created an in-domain vocabulary through Byte-Pair Encoding (BPE) applied to the corpus, leading to a total of 42,384 word pieces. The model utilizes the GPT-2medium as its core network, featuring 24 layers, a hidden size of 1024, and 16 attention heads. This results in a total of 347 million parameters for the BioGPT model.
Galactica is model targeting the parsing of scientific discourse, developing particular tokeneizers for common scientific text (e.g. equations, chemical formulae). It implements the following architectural modifications: GeLU activations [31], 2048-length context window, no biases [32], learned positional embeddings, and 50k-token Byte-Pair Encoding vocabulary [33] from a 2% training data subset. It is available in sizes from 125M to 120B parameters. The largest version that we were able to run on was 1.3B.
Both GPT-4 and ChatGPT builds upon the GPT-3.5 model which is then refined via the Reinforcement Learning with Human Feedback (RLHF) mechanism. Both corpora and human feedback alignment dataset are not reported.
Llama 2 was trained with an extensive corpus of 2 trillion tokens extracted from publicly available sources like Common Crawl (an archive of billions of webpages), Wikipedia, and public domain books from Project Gutenberg. In contrast to the above two models, the training dataset for Llama 2 is ascertainable. Llama 2 exhibits superior performance relative to other open-source language models across diverse external benchmarks, encompassing assessments of logical reasoning, coding proficiency, and general knowledge. However, it is emphasized that this model is just intended to be further trained to tailored to specific requirements. This study evaluates Llama 2 7B, Llama 2 13B and Llama 2 70B.
@@
The final baseline set consists of: GPT-2, GPT-neo (1.3B parameters), BLOOM (3B parameters), BioGPT, BioGPT-large and Galactica via huggingface API [^4] https://huggingface.co; Llama 2 using the llama.cpp [^5] https://github.com/ggerganov/llama.cpp (accessed in October 2023); all above with the default sampling procedure for text generation, i.e. greedy decoding and temperature equal 0; GPT-3 via the OpenAI API; and for ChatGPT and GPT-4 a default OpenAI user interface was used to generate outputs, each prompt in a separate session (accessed in March 2023).
3.3 Framework for the evaluation of Large Language Models on the encoding of factual scientific knowledge
An intrinsic evaluation of the text generated by the models was performed by a human expert according to five criteria: Fluency, Prompt-alignment, Semantic coherence, Factuality and Specificity [34, 35]:
  • Fluency evaluates the quality of the generated text, considering two sub-criteria: syntactic correctness and style. Syntactic correctness assesses whether a sentence conforms to the rules of grammar and expected syntactic structure. Scientific style is determined by the vocabulary of the generated text conforming to a scientific theme (‘reads as a scientific text’).
  • Prompt-alignment refers to the relation of relevance and pragmatic coherence between the prompt and the generated text, whether the generated text is consistent with the input prompt and is related to the entities and context of the prompt.
  • Semantic coherence consists of two levels of analysis: (i) intra-sentence assesses whether the sentence is relating terms surface to expected argument types for a given predicate and (ii) inter-sentence: assessing whether adjacent sentences have a coherent discourse relation [36].
  • Factuality is distinct from the meaning of a sentence which is conveyed by its semantic coherence. A sentence generated by a model is true if there exists a factual statement in the scientific literature, a database, or any other trusted source.
  • Specificity evaluates the alignment between the prompt and the answer regarding their level of abstraction. While some answers can be factually correct, they may not answer to the level of abstraction required by the prompt.
While auto-regressive language models are by design to deliver fluency, prompt-alignment and semantic coherence, text generation during decoding can lead to repetitive, incoherent or meaningless outputs [37]. As they are complementary, the fluency, prompt-alignment and semantic coherence were assessed in one step of the evaluation process (Figure 3, STEP 1). Moreover, the probabilistic nature of the next token prediction task also contrasts with the notion of factuality and large language models are prone to hallucination [38]. Hallucinations are factually incorrect statements which can be difficult to be localised by non-experts due to their expression in a fluent and semantically coherent text. In the context of this study, major hallucinations can be associated with the compound description or relationships with fungi, while minor hallucinations may be associated with the date of first isolation for instance. The factuality of the generated text was assessed in the STEP 2. Importantly, for Relation determination via entity generation, the correctness of the generated entity 2 is verified prior to the verification of the whole description. If incorrect, we assume that the LLM fails to determine the relation. Specificity of the generated text is assessed in the STEP 3. Note that only STEP 1 does not require a domain expert.
Figure 3: The framework for qualitative and efficient evaluation of the Large Language Models in the context of extracting biological relation from these models. The entity 1 stands for chemical compound name. The entity 2 stands for fungus name.
In this study two major tasks were evaluated, described in 2.5. Models were prompted with task-specific prompts (see Table 2) and the generated text was evaluated according to the framework depicted in Figure 3.
@@ 3.4 Text generation tasks
We investigate the ability of LLMs to act as knowledge bases and establish the relation between fungi and antibiotics. Following the evaluation framework proposed in section 2.4, we define two text generation tasks (see the workflow in Figure 2).
Task 1. Chemical compound definition generation
Task 1 uses prompts for chemical compound definition generation, shown in the Table 2 as P1-P7. For the same task we define simple prompts (P1, P2) and context-based prompts, elucidating the type of the subject entity: ‘a compound’ or ‘a substance’ (P4 and P5) [39]. For Galactica, according to authors’ suggestions, prompts were framed as questions, rather than sentences to complete (P3, P6, P7). The entity 1 represents the selected chemical compounds. Performance in Task 1 was calculated for each step of the framework (Fig.3) and is shown in Table 3.
Task 2. Chemical-fungus relation determination via entity generation
Task 2 uses prompts for relation determination via entity generation: {entity 1}- {relation} - {entity 2}, where entity 1 is a chemical compound name and entity 2 is a fungus name (P8-P19, see Table 2). The prompts P13-P18 contain the context at the end, guiding the model towards a relation to a fungus, rather than to another organism capable of producing the compound. Evaluation was performed according to the framework (Fig.3) for the entire task, both with and without context. We report the number of responses meeting the criteria of given STEP in the framework for each prompt (x correct out of 110, see Table 4).
Table 2: Prompts used in Task 1 and Task 2
Task Type of prompts No. Prompt
@@
[…]
context-based prompts P4 {entity 1} is a compound
P5 {entity 1} is a substance
[…]
Task 2: Chemical-fungus relation determination via entity generation simple prompts P8 {entity 1} is isolated from
P9 {entity 1} originally isolated from
P10 {entity 1} is produced by
P11 {entity 1} is a natural chemical compound found in
P12 {entity 1} is a natural product found in
context-based prompts P13 {entity 1} is isolated from the fungus
P14 {entity 1} is isolated from fungi
P15 {entity 1} is isolated from fungi, such as
P16 {entity 1} is produced by the fungus
P17 {entity 1} is produced by fungi
P18 {entity 1} is produced by fungi, such as
[…]
    *Evaluation for Galactica included P3, P6, P7 and P19 prompts only.
@@ 3.5 Expert evaluation
As described in 2.2, we selected ten names of chemical compounds (five chemicals with and five without antibiotic activity). The maximum numbers of tokens to generate, excluding the number of tokens in the prompt is 30, except for ChatGPT, GPT-4 and Llama 2. In total, we investigated outputs for 15 prompts for the models Bloom, GPT-2, GPT-neo, GPT-3, BioGPT, BioGPT-large, ChatGPT, GPT-4 and Llama 2; and four prompts for Galactica. All prompts are presented in Table 2. In both tasks, in order to measure the variance in our small sample size setting, we applied bootstrapping with 1000 iterations, reporting the range (in %) of correct responses for a given prompt and STEP with the 95% confidence interval.
All 1690 outputs generated by the models, i.e. 11 models x 15 prompts x 10 chemical compounds + 1 model (Galactica) x 4 prompts x 10 chemical compounds, were evaluated and labeled according to the proposed framework by 2 reviewers independently of each other, on each of steps 1-3. A total of 1690 records were evaluated on the step 1. 1231 of them were moved to the steps 2 and 3, and then still dually reviewed by two reviewers. During all steps, there were 22 (1.3%) discrepancies, which were resolved by discussion. Two experienced reviewers took part in this systematic analysis. Both reviewers were involved in the study design of the systematic analysis and the definition of the criteria. The reviewers went through a pilot process with the criteria prior to starting the study to ensure they had a similar understanding of the criteria. The evaluation was carried out systematically by two independent experts in the field, one with a PhD in microbiology and chemistry and the other with a PhD in biochemical pharmacology, both having prior experience in the manual assessment and curation of datasets.
@@ 4 Results
4.1 Chemical compound definition generation
This task assesses the ability of the LLM to generate a factual definition of a chemical compound (equal entity 1). For instance with the Prompt = ‘{entity 1} is’ with the entity ‘Atromentin’, a correct answer could have the form ‘Atromentin is a natural product…’. We investigated ten examples of chemical compounds (five with and five without antibiotic activity) for the qualitative evaluation according to the previously described criteria (see section 2.4, Table S.1). As prompts, we provided single entities or entities with the more specific context (see section 2.5). The results are summarized in Table 3. Examples of generated answers with the qualitative evaluation are presented in Table S.2. Note, for ChatGPT, GPT-4 and Llama 2 we did not limit the number of tokens and evaluated the complete output. In total, we manually evaluated 470 outputs in this task.
GPT-2 and GPT-neo: no semantic coherence, no factuality. GPT-2 failed to generate any semantically coherent text that is related to the entity and to the target domain (biomedicine in general). After adding ‘is’ to the prompts, the model generates fluent text only for 2/10 compounds, but with no semantic coherence. Adding context leads to even more fluent outputs but still loses semantic coherence and there is no prompt-alignment. For 7/10 compounds very fluent text is generated that is unrelated to either the associated fungus or the antibiotic activity of the compound.
@@ 4.1 Chemical compound definition generation
GPT-neo, similarly to GPT-2, failed to generate any text or any fluent text. After adding ‘is’ - the model generates (for 8/10) syntactically correct texts unrelated to the given name, resembling to an n-gram scientific fragment unrelated to the mention. After adding context, the model asserts two compounds (Fumitremorgin C, Conidiogenone) as being produced by fungi (Fusarium sphaerospermum, Penicillium notatum, respectively), however under incorrect associations. The model failed to generate factual knowledge.
@@
GPT-3: Semantically coherent answers less sensitive to prompt design. 3/10 factual answers. Hallucinates fungi names from compounds. GPT-3 generated semantically coherent text regardless of the prompt design for all chemical compounds, both with and without context (see example of Atromentin in Table S.2). It generated a similar number of factual answers regardless of the context (30% vs 35%). Interestingly, adding context to the prompt did not increase specificity (‘Ergosterol is … a sterol found in the cell membranes of fungi. It is a major component of the fungal cell membrane, and is similar to cholesterol in animals. It is an important precursor for the synthesis of other compounds, such as ergocalciferol’ vs ‘Ergosterol is a compound … found in the cell membranes of fungi. It is a sterol, a type of lipid molecule, and is the main sterol in the fungal cell membrane. It is similar to cholesterol, which is found in animal cell membranes’). For 2/10 chemicals (Ergosterol and Alternariol) GPT-3 generated factual and specific answers regardless of the prompt and context. GTP-3 generated incorrect fungi names (in 39.3% of incorrect answers) that consists of the compound name in the prompt, e.g. Myriocin - Streptomyces myriocin, Chloromonicilin - Streptomyces chloromyceticus (Table S.3).
@@
Bloom: no factual knowledge, low semantic coherence: biased towards treatments, repeats the same associated terms or statements. For prompts without context, Bloom fails to generate a syntactically correct output. It does not generate any text (for example, ‘Conidiogenone ’), generates the chemical compound entity as a code within a link (for example, ‘org.apache.thrift.TBase("org.apache.thrift.TBase", "flavasperone",’) or generates text with only special characters (for example, ‘Ergosterol ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñ’). The model tends to repeat the same words or phrases for a given entity (for example, ‘Fumitremorgin C is a potent inhibitor of the enzyme, which is responsible for the synthesis of the enzyme, and is therefore a potent inhibitor of the synthesis of the enzyme.’). In addition, the model generates names that does not exist in the literature (for example, ‘Atromentin is a natural product derived from the fruit of the plant Atrium tridentatum.’). When adding the context to the prompts, Bloom is biased towards considering the compounds as a treatment for a variety of diseases. It generates fluent but no semantically coherent text, unrelated or not specific to the compound. Frequently the answer contains repeated statements and facts, e.g. ‘Fumitremorgin C is a substance that is used in the treatment of cancer. It is a natural product that is used in the treatment of cancer. It is a natural product that is’ - despite its factuality, it corresponds to low prompt-alignment and specificity.
@@
BioGPT: low specificity, adding context leads to even lower specificity, leading to the answer being true for any compound. BioGPT generates fluent and syntactically correct text tends to correspond to a more general discourse type (when contrasted to scientific discourse). For prompts ‘{entity 1}’ and ‘{entity 1} is’, only one compound (Ergosterol) has factual answers. For instance, the generated text states that one compound is an antibiotic, but the rest of the sentence is not related to the compound, or the model confuses DNA polymerase alpha/beta/gamma (Table S.2 for Aphidicolin). When more context is provided (prompt with ‘is a compound’), BioGPT generates a factual text with low specificity for 4/10 compounds. It provides a non-specific, universal definition of a compound, which is referred to regardless of the compound name (Table S.2, example: ‘Fumitremorgin C is a compound with a unique structure and a unique mode of action.’ - the same is for Verrucarin A). After adding the context ‘is a substance’, BioGPT recognises that 5 /10 compounds were produced by fungus, but all incorrect.
@@
BioGPT-large: Semantically very coherent. 3/10 factual answers. Hallucinations under high quality, scientific text are difficult to spot. BioGPT-large generated semantically coherent outputs for 80% of the prompts without and 100% prompts with the context. However, the answers were factual only for 30% and 35% for prompts without and with context, respectively. Despite some of them were not specific, BioGPT-large generated answers better fitting to a scientific discourse level, compared to the previous models (Table S.3). Similarly to GPT-3, for 28% of answers with no factuality, BioGPT-large assigned a wrong producer to the compound, e.g. ’Chloromonilicin is a new antibiotic produced by a strain of Streptomyces. It is active against Gram-positive bacteria and fungi.’ (true answer: is produced by Monilinia fructicola). This is of particular concern, as the generated text imitates correct definitions very well, disguising the lack of factual knowledge. Such hallucination is very difficult to spot and a vigilant expert check is required.
Galactica: poor semantic coherence, improves after adding context to the prompt, answers resembles as corpora fragments. 1/10 factual answers. Galactica only for 2/10 compound names can generate text that is related to the target entity and that is aligned to the domain, but in both cases lacks factuality. For one case, the model assigned that the compound was produced by a fungus but misnamed it (‘Aphidicolin is a polyketide antibiotic produced by the fungus Aspergillus nidulans. It is a potent inhibitor of DNA polymerase alpha.’). Adding the context ‘What is the compound {entity 1}?’ improves the output. The text is more meaningful and fluent, and is domain-aligned. However, it still lacks factual knowledge with one exception for Myriocin, which is recognized as an antibiotic. The generated text resembles the fragment from a title of a specific scientific paper than a general description (‘Myriocin, a new antifungal antibiotic. I. Taxonomy, fermentation, isolation, physico-chemical properties and biological activity.’). Moreover, the model recognises that Fumitremorgin C is a mycotoxin but again misnamed the fungus name (‘Fusarium verticillioides’, instead of ‘Aspergillus fumigatus’).
@@
ChatGPT: Fluent, prompt-alignment and semantically coherent descriptions. Hallucinations are difficult to spot. 6/10 factual answers. ChatGPT generated fluent and semantically coherent text for all prompts. For prompts with no context, the statements were longer and divided into paragraphs (mostly three; note, that a limit was not set on the length of generated answer for ChatGPT). The generated text was coherent regardless of the designed prompts (’{entity 1}’ or ’{entity 1} is’), and 6/10 outputs were factual. The other four generated the wrong fungus name or the wrong chemical formula in the answer. Interestingly, the answers often closely resembled Wikipedia-style descriptions (7/10 compounds - Wikipedia, 2/10 - Wikidata). For prompts with the context, the answers were shorter and mostly consisted of one paragraph. For prompts ’{entity 1} is a compound’ the text contained a chemical formula and a description of the chemical structure. However, in 60% of these prompts the chemical formula was incorrect (see Aphidicolin in Table S.2). The answers that were factual had low specificity, and could satisfy the definition of many other chemical compounds. For prompts ’{entity 1} is a substance’, the answer provides additional biological context and informs about the biological activity of the compound or what it is used against. 70% of these answers contained factual knowledge. Comparing to BioGPT-large, ChatGPT produces even more fluent, prompt-alignment and semantically coherent answers. Due to semantic fluency, evaluation of the factual knowledge is very challenging, and requires both the expert knowledge and time-consuming fact verification. In other words, generated answers can be easily regarded as true and factual for a non-expert, creating a real risk of deriving false facts from the model.

### v3

@@ 3.2 Dataset
Investigating the biological knowledge stored in LLMs requires a collection of biological relations that are elicited and referred to in the scientific literature. Research has shown that language models (LM) exhibit bias due to imbalanced distributions in their training corpora, with performance correlated to the frequency of occurrences in the data Henning et al. (2022); Wysocki et al. (2023b); Jung and van der Plas (2024). Scientific publications are part of the training corpus for most LLMs, thus higher prevalence in literature should lead to better performance in factual knowledge extraction. In our study, we created a dataset of biological relations accounting for the amount of supporting evidence (number of associated papers).
Chemical compound-fungus pairs were extracted from the Wikidata Knowledge Graph using the predicate found in taxon (p:P703), along with the number of bibliographic references supporting each relationship (via API). The classification and nomenclature information on the fungi was extracted from the Mycobank database [^2] https://www.mycobank.org. Information about antibiotic activity of chemicals were extracted from the ChEBI ontology [^3] https://www.ebi.ac.uk/chebi/downloadsForward.do.
A subset containing only fungal species producing chemicals with antibiotic activity was first created. The pairs with the most bibliographic references were progressively integrated until obtaining 100 distinct subject fungi, which corresponds to 123 pairs. Then, a complementary subset (with an equal number of pairs) involving produced chemicals but without antibiotic activity was added from the same list of fungi. Again, the pairs with the most bibliographic references were selected. Only molecules with a 3-STAR review in ChEBI (Chemical Entities of Biological Interest) database, indicating the highest level of manual curation and verification by ChEBI curators, were considered and only those that have an annotated role in the antimicrobial agent class (CHEBI:33281) were selected for positive examples of antibiotics. As a result, a dataset with 246 chemical compound-fungus pairs (see Fig. 2, $N$) equally divided between chemicals with and without antibiotic activity was obtained. For the experiments, we randomly selected ten chemicals distributed by antibiotic activity and number of references (five chemicals with and five without antibiotic activity) (Table A.1). These ten chemicals paired with one or more fungi. In total, there were 23 chemical compound-fungus pairs (see Fig. 2, $n_{relations}$) in the analysed subset. The selected ten chemical compounds were defined as entity 1 for two tasks. The fungi were defined as entity 2 in the second task. Of note, the selected 23 chemical-fungus relations are described by multiple scientific publications in PubMed, all published before 2019 (for exact numbers see Table A.1). This ensures these relations are contained in the LLM training corpus, provided that the corpus contains PubMed articles or abstracts.
@@ 3.3 Large Language Models
In this study, we evaluated eleven LLMs (summarised in Table A.2), where six of them were released since July 2022. The models differ in terms of training corpus, vocabulary size, number of parameters, layers and maximum input sequence length.
@@
The final baseline set consists of: GPT-2, GPT-neo (1.3B parameters), BLOOM (3B parameters), BioGPT and BioGPT-large via huggingface API [^4] https://huggingface.co; Llama 2 using the llama.cpp [^5] https://github.com/ggerganov/llama.cpp (accessed in October 2023); all above with the default sampling procedure for text generation, i.e. greedy decoding and temperature equal 0; GPT-3 via the OpenAI API; and for ChatGPT and GPT-4 a default OpenAI user interface was used to generate outputs, each prompt in a separate session (accessed in March 2023). The maximum numbers of tokens to generate, excluding the number of tokens in the prompt is 30, except for ChatGPT, GPT-4 and Llama 2.
@@ 3.4 Text generation tasks
We investigate the ability of LLMs to act as knowledge bases and establish the relation between fungi and antibiotics. Following the framework (3.1), we define two text generation tasks (Fig. 2).
Task 1: Chemical compound definition generation
Task 1 uses prompts for chemical compound definition generation, shown in the Table 1 as P1-P4. For the same task we define simple prompts (P1, P2) and context-based prompts, elucidating the type of the subject entity: ‘a compound’ or ‘a substance’ (P3 and P4) Petroni et al. (2020). The entity 1 represents the selected chemical compounds. We investigated ten examples of chemical compounds for the qualitative evaluation according to the previously described criteria (see section 3.1, Table A.1). Performance in Task 1 in each STEP (Fig. 1) is shown in Table 2. Performance in Task 1 for optimal prompt in each STEP is shown in Table 3. Examples of generated answers with the qualitative evaluation are presented in Table A.3.
Task 2 uses prompts for relation determination via entity generation: {entity 1}- {relation} - {entity 2}, where entity 1 is a chemical compound name and entity 2 is a fungus name (P5-P15, Table 1). The prompts P10-P15 contain the context at the end, guiding the model towards a relation to a fungus, rather than to another organism capable of producing the compound. We investigated ten compounds, the same as in Task 1 described in section 3.2. Evaluation was performed according to the framework (Fig. 1) for the entire task, both with and without context. Performance in Task 2 is each STEP is shown in Table 4. Performance in Task 2 for optimal prompt in each STEP is shown in Table 5.
Table 1: Prompts used in Task 1 and Task 2.
Task Type of prompt No. Prompt
@@
context-based prompts P3 {entity 1} is a compound
P4 {entity 1} is a substance
Task 2: Chemical-fungus relation determination via entity generation simple prompts P5 {entity 1} is isolated from
P6 {entity 1} originally isolated from
P7 {entity 1} is produced by
P8 {entity 1} is a natural chemical compound found in
P9 {entity 1} is a natural product found in
context-based prompts P10 {entity 1} is isolated from the fungus
P11 {entity 1} is isolated from fungi
P12 {entity 1} is isolated from fungi, such as
P13 {entity 1} is produced by the fungus
P14 {entity 1} is produced by fungi
P15 {entity 1} is produced by fungi, such as
@@ 3.5 Expert evaluation
The outputs were evaluated and labeled according to the proposed framework by two experienced reviewers independently of each other, on each of steps 1-3. They are experts in the field, one with a PhD in microbiology and chemistry and the other with a PhD in biochemical pharmacology, both having prior experience in the manual assessment and curation of datasets. Any discrepancies were resolved by discussion. Both reviewers participated in the study design and the definition of criteria. The reviewers went through a pilot process with the criteria prior to starting the study to ensure they had a similar understanding of the criteria.
@@ 4 Results
Evaluation of 11 models, 10 chemical compounds and 15 prompts resulted in a total of 1650 responses: 440 from Task 1 (11x10x4), 1210 from Task 2 (11x10x11). Our framework demonstrated a significant reduction in the number of outputs requiring expert evaluation, decreasing by 33% in Task 1 and 46% in Task 2 (see Fig. 3), by discarding incorrect outputs in STEP 1. For Task 1, only 67% of the outputs proceeded to the factuality check, and ultimately, only 21% required a detailed assessment of specificity. In Task 2, 54% of the outputs needed factuality verification for entity 2, and subsequently, only 21% required comprehensive factuality evaluation of the entire output.
The responses that meet the criteria for each model and for each STEP are reported in Table 2 and 4, and Supp.Fig. A.1, A.2, A.3. Examples of responses generated with qualitative evaluation are presented in Table A.3 and A.4, and optimal prompt with the best performance in Table 3 and 5, and Supp.Fig. A.4, A.5. In the following subsections, we provide detailed results for each individual LLM.
(a) A)
(b) B)
Figure 3: The number of outputs evaluated at each STEP of the framework, showing the percentage reduction compared to the initial number of outputs evaluated in STEP 1 by non-experts. Results are aggregated for 11 models and prompts from the given task: A) Task 1, B) Task 2.
4.1 Task 1: Chemical compound definition generation
@@
GPT-2 and GPT-neo: no semantic coherence, no factuality. GPT-2 failed to generate any semantically coherent text related to the entity and the target domain (biomedicine in general). Adding context leads to more fluent outputs but still loses semantic coherence and there is no prompt-alignment. GPT-neo, similarly to GPT-2, failed to generate any text or any fluent text. After adding context, the model asserts two compounds (Fumitremorgin C, Conidiogenone) as being produced by fungi (Fusarium sphaerospermum, Penicillium notatum, respectively), however, under incorrect associations. The model failed to generate factual knowledge.
@@
GPT-3: Semantically coherent answers less sensitive to prompt design. 40% of factual answers for the best prompt (Table 3). Hallucinates fungi names from compounds. GPT-3 generated semantically coherent text regardless of the prompt design for all chemical compounds, both with and without context (see example of Atromentin in Table A.3). It generated a similar number of factual answers regardless of the context (30% vs 35%). GTP-3 generated incorrect fungi names (in 39.3% of incorrect answers) that consists of the compound name in the prompt, e.g. Myriocin - Streptomyces myriocin, Chloromonicilin - Streptomyces chloromyceticus (Table A.4).
@@
Bloom: no factual knowledge, low semantic coherence: biased towards treatments, repeats the same associated terms or statements. For prompts without context, Bloom fails to generate a syntactically correct output. The model tends to repeat the same words or phrases for a given entity, or generates names that does not exist in the literature. When adding the context to the prompts, Bloom is biased towards considering the compounds as a treatment for a variety of diseases. Frequently the answer contains repeated statements and facts, e.g. ‘Fumitremorgin C is a substance that is used in the treatment of cancer. It is a natural product that is used in the treatment of cancer.’.
@@
BioGPT: low specificity, adding context leads to even lower specificity, leading to a definitional generation for the compound. BioGPT generates fluent and syntactically correct text and tends to correspond to a more general discourse type (when contrasted to scientific discourse). When more context is provided, BioGPT provides a non-specific, universal definition of a compound, which is referred to regardless of the compound name (Table A.3). After adding the prompt context ‘is a substance’, BioGPT recognises that 5/10 compounds were produced by a fungi, but all incorrect.
@@
BioGPT-large: Semantically very coherent. 40% of factual answers for the best prompt (Table 3). Hallucinations under high quality, scientific generated text are difficult to spot. Similarly to GPT-3, for 28% of answers with no factuality, BioGPT-large assigned a wrong producer to the compound, e.g. ’Chloromonilicin is a new antibiotic produced by a strain of Streptomyces. It is active against Gram-positive bacteria and fungi.’ (true answer: is produced by Monilinia fructicola). This is of particular concern, as the generated text imitates correct definitions very well, disguising the lack of factual knowledge. Such hallucination is very difficult to spot and a vigilant expert check is required.
@@
ChatGPT: Fluent, prompt-alignment and semantically coherent descriptions for all prompts. Hallucinations are difficult to spot. 70% of factual answers for the best prompt (Table 3). Interestingly, the answers often closely resembled Wikipedia-style descriptions (7/10 compounds - Wikipedia, 2/10 - Wikidata). For prompts with no context, the statements were longer and divided into paragraphs (mostly three; note, that a limit was not set on the length of generated answer for ChatGPT). Comparing to BioGPT-large, ChatGPT produces even more fluent, prompt-aligned and semantically coherent answers. Due to its semantic fluency, evaluation of the factual knowledge requires both expert knowledge and the time-consuming fact verification. In other words, generated answers can be easily regarded as true and factual for a non-expert, creating a real risk of deriving false facts from the model.